# Wall crossing and the Fourier–Mukai transform for Grassmann flops

### v1

@@ 5.3. Main theorem
${K^{0}_{\mathbb{T}}(X_{-})}$${K^{0}_{\mathbb{T}}(X_{+})}$~${\widetilde{\mathcal{H}}_{X_{-}}}$~${\widetilde{\mathcal{H}}_{X_{+}}.}$$\scriptstyle{\mathbb{FM}}$$\scriptstyle{\Psi_{+}}$$\scriptstyle{\Psi_{-}}$$\scriptstyle{\mathbb{U}}$
@@
Equation (5.1.1) and the analogous statement for $X_{-}$ together with (5.3.2) immediately implies that $\mathbb{U}I_{X_{-}}=\widetilde{I_{X_{+}}}.$ The fact that $\mathbb{U}$ is symplectic and has a well-defined non-equivariant limit is Corollary 3.6. The commuting diagram the follows from Proposition 2.9.
Remark 5.5 (Big $I$-functions and the comparison with [LSW]).
In the special case that $B$ is a point, the wall crossing from $X_{+}$ to $X_{-}$ is the standard Grassmann flop. In this case, by Webb [Web23, Section 6.3], one can obtain big $I$-functions of $X_{\pm}$, denoted $\mathbb{I}_{X_{\pm}}(q_{\pm},\mathbf{x},z)$, as a modification of $I_{X_{\pm}}(q_{\pm},z)$. The respective big $I$-functions generate the entire Lagrangian cones $\mathcal{L}_{X_{\pm}}$, and therefore fully determine the genus-zero Gromov–Witten theory of $X_{\pm}$. A straightforward generalization of the computations above shows that the transformation $\mathbb{U}$ from Theorem 5.4 also identifies these big $I$-functions:
(5.3.3) $\mathbb{U}\mathbb{I}_{X_{-}}(q_{-},\mathbf{x},z)=\widetilde{\mathbb{I}_{X_{+}}}(q_{+},\mathbf{x},z),$
@@
In the case that $B$ is a point, Lutz–Shafi–Webb [LSW] have obtained a different proof that there exists a symplectic transformation, which we will denote here by $\mathbb{U}^{LSW}$, identifying the big $I$-functions of $X_{\pm}$ after the analytic continuation. Rather than comparing to the Fourier–Mukai transform $\mathbb{FM}$ however, the symplectic transformation in $\mathbb{U}^{LSW}$ is obtained directly from wall crossing of the associated abelian GIT quotients $X_{T,\pm}$. Their result is related to Theorem 5.4 as follows. Let $\mathbb{I}_{X_{\pm}}(q_{\pm},\mathbf{x},z)$ be the big $I$-functions of the previous paragraph, and let $g^{*}\mathbb{I}^{\pm}_{G}(\mathbf{y}_{\pm},\mathbf{x},z)$ denote the $I$-functions used in [LSW]. Comparing explicit formulas, we have
(5.3.4) $\mathbb{I}_{X_{\pm}}(q_{\pm},\mathbf{x},z)|_{q_{\pm}=\mathbf{y}_{\pm}}=\mathbb{I}^{\pm}_{G}(\mathbf{y}_{\pm},\mathbf{x},z)|_{\mathbf{q}=1},$
Under the identification $q_{\pm}\mapsto\mathbf{y}_{\pm}$, the path $\delta$ of analytic continuation used in loc. cit. matches that given in Proposition 5.1. In [LSW] the abelian/non-abelian correspondence is used to show that there is a unique linear map $\mathbb{U}^{LSW}$ satisfying
$\mathbb{U}^{LSW}\mathbb{I}^{-}_{G}(\mathbf{y}_{-}\mathbf{x},z)|_{\mathbf{q}=1}=\widetilde{\mathbb{I}^{+}_{G}}(\mathbf{y}_{+}\mathbf{x},z)|_{\mathbf{q}=1}.$
From this, (5.3.4), and (5.3.3), we conclude that $\mathbb{U}=\mathbb{U}^{LSW}$, thus providing a consistency check between the two results.
5.4. Central charge
Motivated by the notion of central charge as given in e.g. [KA23, Iri09], we make a similar definition. The above wall crossing result takes an especially simple form using this language.
@@ 5.4. Central charge
$Z_{X_{\pm}}(E)(Q,q_{\pm},z):=\left\langle I_{X_{\pm}}(Q,q_{\pm},-z),\Psi_{\pm}(E)\right\rangle_{X_{\pm}}.$
@@
For $E\in K^{0}_{\mathbb{T}}(X_{\pm})$, the analytic continuation of $Z_{X_{+}}(\mathbb{FM}(E))(Q,q_{+},z)$ to $q_{+}=\infty$ is $Z_{X_{-}}(E)(Q,q_{-},z)$.
@@
We use the diagram in Theorem 4.2 and the fact that $\mathbb{U}$ preserves the symplectic pairing.
$\widetilde{Z_{X_{+}}}(\mathbb{FM}(E))(Q,q_{+},z)=$ $\left\langle\widetilde{I_{X_{+}}}(Q,q_{+},-z),\Psi_{+}(\mathbb{FM}(E))\right\rangle_{X_{+}}$
$=$ $\left\langle\mathbb{U}I_{X_{-}}(Q,q_{-},-z),\mathbb{U}\Psi_{-}(E)\right\rangle_{X_{+}}$
$=$ $\left\langle I_{X_{-}}(Q,q_{-},-z),\Psi_{-}(E)\right\rangle_{X_{-}}$
$=$ $Z_{X_{-}}(E)(Q,q_{-},z).$
@@
(A.0.1) $\Gamma(1-z)\Gamma(z)=\frac{\pi}{\sin(\pi z)},$
@@
$\frac{\prod_{i=1}^{n}\sin((-z_{i}-y)/2\mathbbm{i}))}{\pi^{n}}\frac{\prod_{i=1}^{n}(-1)^{\beta\cdot M_{i}+e}\Gamma((-z_{i}-y)/2\pi\mathbbm{i}+\beta\cdot M_{i}+e)}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}-y)/2\pi\mathbbm{i}+\beta\cdot L_{j}+e)}.$
@@
(A.0.2) $\sum_{e\geq-\beta\cdot L_{l}}q_{+}^{e+x_{l}/2\pi\mathbbm{i}}\frac{\prod_{i=1}^{n}\sin((-z_{i}+x_{l})/2\mathbbm{i})}{\pi^{n}}\frac{\prod_{i=1}^{n}(-1)^{\beta\cdot M_{i}+e}\Gamma((-z_{i}+x_{l})/2\pi\mathbbm{i}+\beta\cdot M_{i}+e)}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}+x_{l})/2\pi\mathbbm{i}+\beta\cdot L_{j}+e)}$
@@
$\left[\prod_{j=1}^{n}(\partial_{q}-x_{j}/2\pi\mathbbm{i}+\beta\cdot L_{j})-q_{+}(-1)^{n}\prod_{j=1}^{n}(\partial_{q}-z_{j}/2\pi\mathbbm{i}+\beta\cdot M_{j})\right]\Phi(q_{+})=0,$
where $\partial_{q}=q_{+}\frac{\partial}{\partial q_{+}}$. This differential equation has singularities at $0,\infty,$ and $(-1)^{n}$. Following [CIJ18], our path $\gamma$ of analytic continuation in the $\log(q_{+})$ will go through $\mathfrak{Im}(\log(q_{+}))=(n-1)\pi\mathbbm{i}$ when $\mathfrak{R}(\log(q_{+}))$ is close to 0. See Figure 1 for a picture of the path. We will use the Mellin–Barnes method to compute this analytic continuation of $H_{\beta,l}^{+}$ to $q_{+}=\infty$.
We may rewrite (A.0.2) as a contour integral
(A.0.3) $\frac{1}{2\pi\mathbbm{i}}\frac{\prod_{i=1}^{n}\sin((-z_{i}+x_{l})/2\mathbbm{i})}{\pi^{n}}\int_{C}\Gamma(s)\Gamma(1-s)q_{+}^{s+x_{l}/2\pi\mathbbm{i}}\cdot$
$e^{-\pi\mathbbm{i}(\sum_{i=1}^{n}\beta\cdot M_{i}+(n-1)\cdot s)}\frac{\prod_{i=1}^{n}\Gamma((-z_{i}+x_{l})/2\pi\mathbbm{i}+\beta\cdot M_{i}+s)}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}+x_{l})/2\pi\mathbbm{i}+\beta\cdot L_{j}+s)},$
@@
$-(-z_{k}+x_{l})/2\pi\mathbbm{i}-d\hskip 14.22636pt1\leq k\leq n,d\in\mathbb{Z}_{\geq\beta\cdot M_{k}}.$
@@
(A.0.4) $-\sum_{1\leq k\leq n}\sum_{d\geq\beta\cdot M_{k}}\Gamma(-(-z_{k}+x_{l})/2\pi\mathbbm{i}-d)\Gamma(1+(-z_{k}+x_{l})/2\pi\mathbbm{i}+d)q_{+}^{z_{k}/2\pi\mathbbm{i}-d}$
$\cdot$ $\frac{e^{-\pi\mathbbm{i}(\sum_{i=1}^{n}\beta\cdot M_{i}+(n-1)\cdot(-(-z_{k}+x_{l})/2\pi\mathbbm{i}-d)+d-\beta\cdot M_{k})}}{(d-\beta\cdot M_{k})!}\frac{\prod_{i\neq k}\Gamma((-z_{i}+z_{k})/2\pi\mathbbm{i}+\beta\cdot M_{i}-d)}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}+z_{k})/2\pi\mathbbm{i}+\beta\cdot L_{j}-d)}$
$\cdot$ $\frac{\prod_{i=1}^{n}\sin((-z_{i}+x_{l})/2\mathbbm{i})}{\pi^{n}}$
$=$ $-\sum_{1\leq k\leq n}\sum_{d\geq\beta\cdot M_{k}}\Gamma(-(-z_{k}+x_{l})/2\pi\mathbbm{i})\Gamma(1+(-z_{k}+x_{l})/2\pi\mathbbm{i})q_{+}^{z_{k}/2\pi\mathbbm{i}-d}$
$\cdot$ $\frac{e^{(n-1)(x_{l}-z_{k})/2}}{(d-\beta\cdot M_{k})!\prod_{j=1}^{n}\Gamma(1+(-x_{j}+z_{k})/2\pi\mathbbm{i}+\beta\cdot L_{j}-d)\prod_{i\neq k}\Gamma(1-(-z_{i}+z_{k})/2\pi\mathbbm{i}-\beta\cdot M_{i}+d)}$
$\cdot$ $\frac{\prod_{i=1}^{n}\sin((-z_{i}+x_{l})/2\mathbbm{i})}{\pi\prod_{i\neq k}\sin((-z_{i}+z_{k})/2\mathbbm{i})}$
$=$ $\sum_{1\leq k\leq n}\sum_{d\geq\beta\cdot M_{k}}e^{(n-1)(x_{l}-z_{k})/2}\frac{\prod_{i\neq k}\sin((-z_{i}+x_{l})/2\mathbbm{i})}{\prod_{i\neq k}\sin((-z_{i}+z_{k})/2\mathbbm{i})}$
$\cdot$ $\frac{q_{+}^{z_{k}/2\pi\mathbbm{i}-d}}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}+z_{k})/2\pi\mathbbm{i}+\beta\cdot L_{j}-d)\prod_{i=1}^{n}\Gamma(1-(-z_{i}+z_{k})/2\pi\mathbbm{i}-\beta\cdot M_{i}+d)}$
where we have again simplified using the Gamma function identity (A.0.1).
Let $B_{k}^{-}$ denote the $k$th fixed locus of $X_{-}$. By (4.1.3), observe that $H_{\beta,k}^{-}:=H_{X_{-},\beta}|_{B_{k}^{-}}$ is given by
$\sum_{d\geq\beta\cdot M_{k}}\frac{q_{-}^{-z_{k}/2\pi\mathbbm{i}+d}}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}+z_{k})/2\pi\mathbbm{i}+\beta\cdot L_{j}-d)\Gamma(1+(z_{j}-z_{k})/2\pi\mathbbm{i}-\beta\cdot M_{j}+d)}.$
The result then follows immediately from (4.1.5).

### v2

@@ 5.3 Main theorem
The linear transformation $\mathbb{U}$ is symplectic, has a well-defined non-equivariant limit, and satisfies
@@
${K^{0}_{\mathbb{T}}(X_{-})}$${K^{0}_{\mathbb{T}}(X_{+})}$~${\widetilde{\mathcal{H}}_{X_{-}}}$~${\widetilde{\mathcal{H}}_{X_{+}}.}$$\scriptstyle{\mathbb{FM}}$$\scriptstyle{\Psi_{-}}$$\scriptstyle{\Psi_{+}}$$\scriptstyle{\mathbb{U}}$
@@
Equation (5.1) and the analogous statement for $X_{-}$ together with (5.13) immediately implies that $\mathbb{U}I_{X_{-}}=\widetilde{I_{X_{+}}}.$ The fact that $\mathbb{U}$ is symplectic and has a well-defined non-equivariant limit is Corollary 3.6. The commuting diagram the follows from Proposition 2.10. ∎
Remark 5.5 (big $I$-functions and the comparison with [28]).
In the special case that $B$ is a point, the wall crossing from $X_{+}$ to $X_{-}$ is the standard Grassmann flop. In this case, by Webb [33, Section 6.3], one can obtain big $I$-functions of $X_{\pm}$, denoted $\mathbb{I}_{X_{\pm}}(q_{\pm},\mathbf{x},z)$, as a modification of $I_{X_{\pm}}(q_{\pm},z)$. The respective big $I$-functions generate the entire Lagrangian cones $\mathcal{L}_{X_{\pm}}$, and therefore fully determine the genus-zero Gromov–Witten theory of $X_{\pm}$. A straightforward generalization of the computations above shows that the transformation $\mathbb{U}$ from Theorem 5.4 also identifies these big $I$-functions:
$\mathbb{U}\mathbb{I}_{X_{-}}(q_{-},\mathbf{x},z)=\widetilde{\mathbb{I}_{X_{+}}}(q_{+},\mathbf{x},z),$ (5.14)
@@
In the case that $B$ is a point, Lutz–Shafi–Webb [28] have obtained a different proof that there exists a symplectic transformation, which we will denote here by $\mathbb{U}^{\rm LSW}$, identifying the big $I$-functions of $X_{\pm}$ after the analytic continuation. Rather than comparing to the Fourier–Mukai transform $\mathbb{FM}$ however, the symplectic transformation in $\mathbb{U}^{\rm LSW}$ is obtained directly from wall crossing of the associated abelian GIT quotients $X_{T,\pm}$. Their result is related to Theorem 5.4 as follows. Let $\mathbb{I}_{X_{\pm}}(q_{\pm},\mathbf{x},z)$ be the big $I$-functions of the previous paragraph, and let $g^{*}\mathbb{I}^{\pm}_{G}(\mathbf{y}_{\pm},\mathbf{x},z)$ denote the $I$-functions used in [28]. Comparing explicit formulas, we have
$\mathbb{I}_{X_{\pm}}(q_{\pm},\mathbf{x},z)|_{q_{\pm}=\mathbf{y}_{\pm}}=\mathbb{I}^{\pm}_{G}(\mathbf{y}_{\pm},\mathbf{x},z)|_{\mathbf{q}=1}.$ (5.15)
Under the identification $q_{\pm}\mapsto\mathbf{y}_{\pm}$, the path $\delta$ of analytic continuation used in loc. cit. matches that given in Proposition 5.1. In [28], the abelian/non-abelian correspondence is used to show that there is a unique linear map $\mathbb{U}^{\rm LSW}$ satisfying
$\mathbb{U}^{\rm LSW}\mathbb{I}^{-}_{G}(\mathbf{y}_{-}\mathbf{x},z)|_{\mathbf{q}=1}=\widetilde{\mathbb{I}^{+}_{G}}(\mathbf{y}_{+}\mathbf{x},z)|_{\mathbf{q}=1}.$
From this, (5.15) and (5.14), we conclude that $\mathbb{U}=\mathbb{U}^{\rm LSW}$, thus providing a consistency check between the two results.
5.4 Central charge
Motivated by the notion of central charge as given, e.g., in [1, 22], we make a similar definition. The above wall crossing result takes an especially simple form using this language.
@@ 5.4 Central charge
$Z_{X_{\pm}}(E)(Q,q_{\pm},z):=\langle I_{X_{\pm}}(Q,q_{\pm},-z),\Psi_{\pm}(E)\rangle_{X_{\pm}}.$
@@
For $E\in K^{0}_{\mathbb{T}}(X_{-})$, the analytic continuation of $Z_{X_{+}}(\mathbb{FM}(E))(Q,q_{+},z)$ to ${q_{+}\!=\!\infty}$ is $Z_{X_{-}}(E)(Q,q_{-},z)$.
@@
We use the diagram in Theorem 4.2 and the fact that $\mathbb{U}$ preserves the symplectic pairing,
$\widetilde{Z_{X_{+}}}(\mathbb{FM}(E))(Q,q_{+},z)$ $=\bigl{\langle}\widetilde{I_{X_{+}}}(Q,q_{+},-z),\Psi_{+}(\mathbb{FM}(E))\bigr{\rangle}_{X_{+}}$
$=\langle\mathbb{U}I_{X_{-}}(Q,q_{-},-z),\mathbb{U}\Psi_{-}(E)\rangle_{X_{+}}$
$=\langle I_{X_{-}}(Q,q_{-},-z),\Psi_{-}(E)\rangle_{X_{-}}=Z_{X_{-}}(E)(Q,q_{-},z).$
@@
$\Gamma(1-z)\Gamma(z)=\frac{\pi}{\sin(\pi z)},$ (A.1)
@@
$\frac{\prod_{i=1}^{n}\sin((-z_{i}-y)/2{\rm i}))}{\pi^{n}}\frac{\prod_{i=1}^{n}(-1)^{\beta\cdot M_{i}+e}\Gamma((-z_{i}-y)/2\pi{\rm i}+\beta\cdot M_{i}+e)}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}-y)/2\pi{\rm i}+\beta\cdot L_{j}+e)}.$
@@
$\sum_{e\geq-\beta\cdot L_{l}}q_{+}^{e+x_{l}/2\pi{\rm i}}$
[…]
@@
$\left[\prod_{j=1}^{n}(\partial_{q}-x_{j}/2\pi{\rm i}+\beta\cdot L_{j})-q_{+}(-1)^{n}\prod_{j=1}^{n}(\partial_{q}-z_{j}/2\pi{\rm i}+\beta\cdot M_{j})\right]\Phi(q_{+})=0,$
where $\partial_{q}=q_{+}\frac{\partial}{\partial q_{+}}$. This differential equation has singularities at $0,\infty,$ and $(-1)^{n}$. Following [14], our path $\gamma$ of analytic continuation in the $\log(q_{+})$ will go through $\mathfrak{Im}(\log(q_{+}))=(n-1)\pi{\rm i}$ when $\mathfrak{Re}(\log(q_{+}))$ is close to 0. See Figure 1 for a picture of the path. We will use the Mellin–Barnes method to compute this analytic continuation of $H_{\beta,l}^{+}$ to $q_{+}=\infty$.
We may rewrite (A.2) as a contour integral
$\frac{1}{2\pi{\rm i}}\frac{\prod_{i=1}^{n}\sin((-z_{i}+x_{l})/2{\rm i})}{\pi^{n}}$ (A.3)
×$\times\!\int_{C}\Gamma(s)\Gamma(1-s)q_{+}^{s+x_{l}/2\pi{\rm i}}{\rm e}^{-\pi{\rm i}(\sum_{i=1}^{n}\beta\cdot M_{i}+(n-1)\cdot s)}\frac{\prod_{i=1}^{n}\Gamma((-z_{i}+x_{l})/2\pi{\rm i}+\beta\cdot M_{i}+s)}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}+x_{l})/2\pi{\rm i}+\beta\cdot L_{j}+s)},$
@@
$-(-z_{k}+x_{l})/2\pi{\rm i}-d,\qquad 1\leq k\leq n,d\in\mathbb{Z}_{\geq\beta\cdot M_{k}}.$
@@
$-\sum_{1\leq k\leq n}\sum_{d\geq\beta\cdot M_{k}}\Gamma(-(-z_{k}+x_{l})/2\pi{\rm i}-d)\Gamma(1+(-z_{k}+x_{l})/2\pi{\rm i}+d)q_{+}^{z_{k}/2\pi{\rm i}-d}$
×$\times\frac{{\rm e}^{-\pi{\rm i}(\sum_{i=1}^{n}\beta\cdot M_{i}+(n-1)\cdot(-(-z_{k}+x_{l})/2\pi{\rm i}-d)+d-\beta\cdot M_{k})}}{(d-\beta\cdot M_{k})!}\frac{\prod_{i\neq k}\Gamma((-z_{i}+z_{k})/2\pi{\rm i}+\beta\cdot M_{i}-d)}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}+z_{k})/2\pi{\rm i}+\beta\cdot L_{j}-d)}$
×$\times\frac{\prod_{i=1}^{n}\sin((-z_{i}+x_{l})/2{\rm i})}{\pi^{n}}$
$=-\sum_{1\leq k\leq n}\sum_{d\geq\beta\cdot M_{k}}\Gamma(-(-z_{k}+x_{l})/2\pi{\rm i})\Gamma(1+(-z_{k}+x_{l})/2\pi{\rm i})q_{+}^{z_{k}/2\pi{\rm i}-d}$
×$\times\frac{{\rm e}^{(n-1)(x_{l}-z_{k})/2}}{(d-\beta\cdot M_{k})!\prod_{j=1}^{n}\Gamma(1+(-x_{j}+z_{k})/2\pi{\rm i}+\beta\cdot L_{j}-d)}$
×$\times\frac{\prod_{i=1}^{n}\sin((-z_{i}+x_{l})/2{\rm i})}{\prod_{i\neq k}\Gamma(1-(-z_{i}+z_{k})/2\pi{\rm i}-\beta\cdot M_{i}+d)\cdot\pi\prod_{i\neq k}\sin((-z_{i}+z_{k})/2{\rm i})}$
$=\sum_{1\leq k\leq n}\sum_{d\geq\beta\cdot M_{k}}{\rm e}^{(n-1)(x_{l}-z_{k})/2}\frac{\prod_{i\neq k}\sin((-z_{i}+x_{l})/2{\rm i})}{\prod_{i\neq k}\sin((-z_{i}+z_{k})/2{\rm i})}$
×$\times\frac{q_{+}^{z_{k}/2\pi{\rm i}-d}}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}+z_{k})/2\pi{\rm i}+\beta\cdot L_{j}-d)\prod_{i=1}^{n}\Gamma(1-(-z_{i}+z_{k})/2\pi{\rm i}-\beta\cdot M_{i}+d)},$
where we have again simplified using the Gamma function identity (A.1).
Let $B_{k}^{-}$ denote the $k$th fixed locus of $X_{-}$. By (4.1), observe that $H_{\beta,k}^{-}:=H_{X_{-},\beta}|_{B_{k}^{-}}$ is given by
$\sum_{d\geq\beta\cdot M_{k}}\frac{q_{-}^{-z_{k}/2\pi{\rm i}+d}}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}+z_{k})/2\pi{\rm i}+\beta\cdot L_{j}-d)\Gamma(1+(z_{j}-z_{k})/2\pi{\rm i}-\beta\cdot M_{j}+d)}.$
The result then follows immediately from (4.3).
Acknowledgements
We would like to thank Tom Coates, Wendelin Lutz, Jeongseok Oh, Yongbin Ruan, Ed Segal, Qaasim Shafi, and Rachel Webb for the useful discussions. We also thank the anonymous referees for valuable comments and suggestions on a previous draft. N. Priddis is supported by the Simons Foundation Travel Grant 586691. M. Shoemaker is supported by the Simons Foundation Travel Grant 958189. Y. Wen is supported by a KIAS Individual Grant (MG083902) at the Korea Institute for Advanced Study.